# Investigation about a statement equivalent to Riemann Hypothesis (RH) applied to Dirichlet primitive L functions

### v15

@@ 3.4. LEMMA 5: Euler Product on critical line
As $p_{max}$ grows, $\rho^{+}(\epsilon,p_{max})\rightarrow\rho^{-}(\epsilon,p_{max})$ for suitable small $\epsilon>0$, both approach 1, but, from below. See (3.47) and ( 3.48). So $X(\epsilon>0,p_{max})=\rho^{\pm}(\epsilon,p_{max})X(\epsilon=0,p_{max})$ for $p_{max}$ big enough. Then we can affirm that (3.50) is zero, and, identity (3.49) forces $X(\epsilon>0,p_{max})=-\left.\frac{\Delta\angle[L_{EP}(t,\epsilon>0,\chi)]}{\Delta t}\right|_{p^{*}\ fixed,\ p_{max}}$ to reach from below $\left.\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi)]}{\Delta t}\right|_{p^{*}\ fixed,\ p_{max}}$ as $\epsilon\rightarrow 0^{+}$ for big $p_{max}$. This is (3.23), that holds $\forall p^{*}$ big. So we can state :
Corollary of Lemma 6. From (3.46 ) and from (3.13) we can write for $\epsilon\rightarrow 0^{+}$:
@@
PROOF. The difference in square brackets is between mean values in $t\pm\frac{\Delta t}{2}$ (3.5) of same function with different $\epsilon$, that in $|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}\ ,\ \epsilon\rightarrow 0$ is smooth in $t$, (i.e. $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$ for $\epsilon=0$), so, for $p^{*}$ big the difference of the mean in the intervals (3.5) is close to the function on the right in (3.51). Besides (3.23) holds for every big $p^{*}$, so, we can work on (3.46 ). For [2, p. 186 (Cauchy condition for series)] : a series $\sum a_{n}=X(\epsilon,p_{max})$ converges iff $\forall\delta>0\ \exists P:\forall n>P$ we have: $|\sum_{n>P}^{n=p+P}a_{n}|<\delta\ \ \forall p\geq 1$ ( i.e. the difference to the limit is $\delta$).
@@
we can find a certain $\epsilon^{\prime}$ and $p^{\prime}_{max}$: $|X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)-X(\epsilon^{\prime}>0,p_{max}\rightarrow\infty)|<\delta$,and, $|X(\epsilon=0,p^{\prime}_{max}<\infty)-X(\epsilon=0,p_{max}\rightarrow\infty)|<\delta$. So we must have : $X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)-X(\epsilon=0,p^{\prime}_{max}<\infty)|>3\delta$ , i.e : $X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)>X(\epsilon=0,p^{\prime}_{max}<\infty)$, i.e. $\rho^{\pm}(\epsilon^{\prime},p^{\prime}_{max})>1$. But this is not possible for $p^{\prime}_{max}$ big, and $\forall p^{*}$, see (3.45 ), (3.47). See Lemma 6 (3.23). So (3.51) must be $\geq 0$.
@@
In [4, p. 6] is reported a statement as equivalent to R.H. for $\zeta(s)$: ”The Riemann hypothesis is equivalent to the statement that all local maxima of $\xi(t,\epsilon=0)$ are positive and all local minima are negative.”. Let us see why. Suppose we are close to extremal points. Then $\Re[\xi(s)]$, is an harmonic function, obeying to Laplace equation $\frac{\partial^{2}\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon^{2}}+\frac{\partial^{2}\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t^{2}}=0$, with null total curvature. See (2.8). As $\frac{\partial\Im[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t}$, and $\Im[\xi(\frac{1}{2}+it)]=0$, then extremal points on critical line (that are saddle points) are first candidate to look for off-critical line zeros. So, at extremal points (relative maxima end minima), if the curvature is toward $\epsilon,t]$ plane along $t$, it will be in the opposite direction along $\epsilon$. Simmetrically if the curvature is toward $\epsilon,t]$ plane along $\epsilon$, it will be in the opposite direction along $t$. Only in second case off-critical line zeros are possible because we have for example $\Re[\xi(t_{0},\epsilon_{0}=0)]>0$ and $\Re[\xi(t,\epsilon_{0}+\Delta\epsilon>0)]<\Re[\xi(t_{0},\epsilon_{0}=0)]$, i.e. we are tending toward the plane $\epsilon,t]$ along $\epsilon$. Instead in first case we have for example $\Re[\xi(t_{0},\epsilon_{0}=0)]>0$ and $\Re[\xi(t,\epsilon_{0}+\Delta\epsilon>0)]>\Re[\xi(t_{0},\epsilon_{0}=0)]$, i.e. we are increasing distance from plane $\epsilon,t]$ along $\epsilon$. In first case the curvature along $t$ must be negative, i.e. $\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$ so $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$. In second case we have instead $\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}>0$ so $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}>0$. Notice the sign of the product $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$ is invariant also if we choose $\Re[\xi(t_{0},\epsilon_{0}=0)]<0$ if all local maxima of $\xi(t,\epsilon=0)$ are positive and all local minima are negative. This means that if all extremal points of $\Re[\xi(t_{0},\epsilon_{0}=0)]$ comply with this rule, no off-critical line zero is possible because $\frac{\partial\Im[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t}\ \$. So $|\Im[\xi(\frac{1}{2}+\epsilon+it)]|$ grows with $\epsilon$ far from $\Re[\xi(\frac{1}{2}+it)]$ extremal points. We do not say that the set of points of $(t,\epsilon)$ plane where $\Im[\xi(s)]=0$ are perpendicular to critical line at $t_{0}$ where $\Re[\xi(\frac{1}{2}+it_{0})]$ is extremal, but, the truth is not far from it [12, p. 3] , and [13, p. 9,fig 6].
@@ 4. Toward the proof $\forall t$ and $\forall$ modulus $q$ of primitive (odd) characters
Here we find same results of appendix A without Stirling formula like in expression (A.5). The purpose is to evaluate (A.13) in the $t=0$ neighborhood. The starting point is the formula ([7, p. 8]) )
@@
Without Stirling formula as in (A.14 ) , the equivalent relation of $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]$ in (3.22 ) or (3.21), is (4.4). See fig. (2), and fig.(14) for respectively odd and even characters and fig 15 for $\zeta(s)(s-1)$ .
@@ 4.1. Odd characters ($\alpha=1$), $\forall t$, and $\forall q$
can be used also locally if the zeros happen at $t$ values in which same situation (4.5) is locally met. The zeros of $L(s,\chi_{primitive})$ distribute like the peaks of the expression (3.13 ) that we know occurs exactly at the zero of $L(s,\chi_{primitive})$. Irrespective if RH is verified or not.
@@
Table 1. Phase of characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=5$. In column $n=0$ the modulus ( we mean : $\sqrt{\Re^{2}[\chi]+\Im^{2}[\chi]}$ ) is zero. Elsewhere it is 1. When $q$ is prime only principal and primitive characters are present. See [3, p. 168].Notice in each, but the first, row $\sum_{h<q:gcd(h,q)=1}e^{i\ \angle[\chi_{h}]}=0$ for $\chi$ not principal see [3, p. 256]. Table loaded from [11] and adapted for (3.13 ) .
@@
Figure 3. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in not principal character case $q=5$ ( see tab. 1) , with $p_{max}=p^{*}=158\times 10^{6}$ ( here are used the first 8868881 primes in (3.13 ) ).In wolfram demostration project - Dirichlet L-Functions and their zeros we can compute, by zeroing real and imaginary parts, the zeros closer to $t=0$ of what here are called X1 and X3, the result is: $\pm 9.443$, $\pm 8.457$, $\pm 6.184,\ \ \pm 4.133\$. Very comparable with results in figure.
Figure 4. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=5$, with $p_{max}=(p^{*})^{3}=45.8\times 10^{6}$ ( here are used the first 2763823 primes in (3.13 ) ).The peaks are lower and fatter with respect to fig. 3, but, the inter-peaks behavior is much more close to $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$, or, better, to (4.4), that follows different behavior close to $t=0$ if $\alpha=0$ (even chrs) or $\alpha=1$ (odd chrs).
Figure 5. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=3$ and $q=4$, with $p_{max}=(p^{*})^{3}=158\times 10^{6}$ ( here are used the first 8868881 primes).For $q=3$, and, $\chi_{1}(n)=(0,1,-1)$ or $\angle[\chi_{1}(n)]=(undefined,0,\pi)$ to comply with format of tab. 1. For $q=4$, and $\chi_{1}(n)=(0,1,0,-1)$, or, $\angle[\chi_{1}(n)]=(undefined,0,undefined,\pi)$ to comply with format of tab. 1. Notice close to $t=0$ the characteristic behavior of odd primitive L-functions. Compare with fig. 2. Until now the only example of even primitive L-function is phase variation of X2Q5 in fig. 3, and fig. 4. In this case compare with fig. 14.
@@
Table 4. Characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=9$, $w=e^{i\pi/3}$. I Table loaded from [11] .
Figure 6. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab. 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes ).
Figure 7. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab. 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes).
Figure 8. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) flipped vertically. So Lemma 2 is verified.
Figure 9. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes). Notice that inter-peaks level is the same of fig, (14) flipped vertically. So Lemma 2 is verified.
Figure 10. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 11. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 12. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (14) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 13. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (14) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.

### v16

@@ 3.4. LEMMA 5: Euler Product on critical line
As $p_{max}$ grows, $\rho^{+}(\epsilon,p_{max})\rightarrow\rho^{-}(\epsilon,p_{max})$ for suitable small $\epsilon>0$, both approach 1, but, from below. See (3.50) and ( 3.51). So $X(\epsilon>0,p_{max})=\rho^{\pm}(\epsilon,p_{max})X(\epsilon=0,p_{max})$ for $p_{max}$ big enough. Then we can affirm that (3.53) is zero, and, identity (3.52) forces $X(\epsilon>0,p_{max})=-\left.\frac{\Delta\angle[L_{EP}(t,\epsilon>0,\chi)]}{\Delta t}\right|_{p^{*}\ fixed,\ p_{max}}$ to reach from below $\left.\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi)]}{\Delta t}\right|_{p^{*}\ fixed,\ p_{max}}$ as $\epsilon\rightarrow 0^{+}$ for big $p_{max}$. This is (3.26), that holds $\forall p^{*}$ big. So we can state :
Corollary of Lemma 6. From (3.49 ) and from (3.16) we can write for $\epsilon\rightarrow 0^{+}$:
@@
PROOF. The difference in square brackets is between mean values in $t\pm\frac{\Delta t}{2}$ (3.6) of same function with different $\epsilon$, that in $|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}\ ,\ \epsilon\rightarrow 0$ is smooth in $t$, (i.e. $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$ for $\epsilon=0$), so, for $p^{*}$ big the difference of the mean in the intervals (3.6) is close to the function on the right in (3.54). Besides (3.26) holds for every big $p^{*}$, so, we can work on (3.49 ). For [2, p. 186 (Cauchy condition for series)] : a series $\sum a_{n}=X(\epsilon,p_{max})$ converges iff $\forall\delta>0\ \exists P:\forall n>P$ we have: $|\sum_{n>P}^{n=p+P}a_{n}|<\delta\ \ \forall p\geq 1$ ( i.e. the difference to the limit is $\delta$).
@@
we can find a certain $\epsilon^{\prime}$ and $p^{\prime}_{max}$: $|X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)-X(\epsilon^{\prime}>0,p_{max}\rightarrow\infty)|<\delta$,and, $|X(\epsilon=0,p^{\prime}_{max}<\infty)-X(\epsilon=0,p_{max}\rightarrow\infty)|<\delta$. So we must have : $X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)-X(\epsilon=0,p^{\prime}_{max}<\infty)|>3\delta$ , i.e : $X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)>X(\epsilon=0,p^{\prime}_{max}<\infty)$, i.e. $\rho^{\pm}(\epsilon^{\prime},p^{\prime}_{max})>1$. But this is not possible for $p^{\prime}_{max}$ big, and $\forall p^{*}$, see (3.48 ), (3.50). See Lemma 6 (3.26). So (3.54) must be $\geq 0$.
@@
In [4, p. 6] is reported a statement as equivalent to R.H. for $\zeta(s)$: ”The Riemann hypothesis is equivalent to the statement that all local maxima of $\xi(t,\epsilon=0)$ are positive and all local minima are negative.”. Let us see why. Suppose we are close to extremal points. Then $\Re[\xi(s)]$, is an harmonic function, obeying to Laplace equation $\frac{\partial^{2}\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon^{2}}+\frac{\partial^{2}\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t^{2}}=0$, with null total curvature. See (2.8). As $\frac{\partial\Im[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t}$, and $\Im[\xi(\frac{1}{2}+it)]=0$, then extremal points on critical line (that are saddle points) are first candidate to look for off-critical line zeros. So, at extremal points (relative maxima end minima), if the curvature is toward $\epsilon,t]$ plane along $t$, it will be in the opposite direction along $\epsilon$. Simmetrically if the curvature is toward $\epsilon,t]$ plane along $\epsilon$, it will be in the opposite direction along $t$. Only in second case off-critical line zeros are possible because we have for example $\Re[\xi(t_{0},\epsilon_{0}=0)]>0$ and $\Re[\xi(t,\epsilon_{0}+\Delta\epsilon>0)]<\Re[\xi(t_{0},\epsilon_{0}=0)]$, i.e. we are tending toward the plane $\epsilon,t]$ along $\epsilon$. Instead in first case we have for example $\Re[\xi(t_{0},\epsilon_{0}=0)]>0$ and $\Re[\xi(t,\epsilon_{0}+\Delta\epsilon>0)]>\Re[\xi(t_{0},\epsilon_{0}=0)]$, i.e. we are increasing distance from plane $\epsilon,t]$ along $\epsilon$. In first case the curvature along $t$ must be negative, i.e. $\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$ so $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$. In second case we have instead $\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}>0$ so $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}>0$. Notice the sign of the product $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$ is invariant also if we choose $\Re[\xi(t_{0},\epsilon_{0}=0)]<0$ if all local maxima of $\xi(t,\epsilon=0)$ are positive and all local minima are negative. This means that if all extremal points of $\Re[\xi(t_{0},\epsilon_{0}=0)]$ comply with this rule, no off-critical line zero is possible because $\frac{\partial\Im[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t}\ \$. So $|\Im[\xi(\frac{1}{2}+\epsilon+it)]|$ grows with $\epsilon$ far from $\Re[\xi(\frac{1}{2}+it)]$ extremal points. We do not say that the set of points of $(t,\epsilon)$ plane where $\Im[\xi(s)]=0$ are perpendicular to critical line at $t_{0}$ where $\Re[\xi(\frac{1}{2}+it_{0})]$ is extremal, but, the truth is not far from it [13, p. 3] , and [14, p. 9,fig 6].
@@ 4. Toward the proof $\forall t$ and $\forall$ modulus $q$ of primitive (odd) characters
Here we find same results of appendix A without Stirling formula like in expression (A.5). The purpose is to evaluate (A.13) in the $t=0$ neighborhood. The starting point is the formula ([8, p. 8]) )
@@
Without Stirling formula as in (A.14 ) , the equivalent relation of $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]$ in (3.25 ) or (3.24), is (4.4). See fig. (2), and fig.(15) for respectively odd and even characters and fig 16 for $\zeta(s)(s-1)$ .
@@ 4.1. Odd characters ($\alpha=1$), $\forall t$, and $\forall q$
can be used also locally if the zeros happen at $t$ values in which same situation (4.5) is locally met. The zeros of $L(s,\chi_{primitive})$ distribute like the peaks of the expression (3.16 ) that we know occurs exactly at the zero of $L(s,\chi_{primitive})$. Irrespective if RH is verified or not.
@@
Table 1. Phase of characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=5$. In column $n=0$ the modulus ( we mean : $\sqrt{\Re^{2}[\chi]+\Im^{2}[\chi]}$ ) is zero. Elsewhere it is 1. When $q$ is prime only principal and primitive characters are present. See [3, p. 168].Notice in each, but the first, row $\sum_{h<q:gcd(h,q)=1}e^{i\ \angle[\chi_{h}]}=0$ for $\chi$ not principal see [3, p. 256]. Table loaded from [12] and adapted for (3.16 ) .
@@
Figure 3. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in not principal character case $q=5$ ( see tab. 1) , with $p_{max}=p^{*}=158\times 10^{6}$ ( here are used the first 8868881 primes in (3.16 ) ).In wolfram demostration project - Dirichlet L-Functions and their zeros we can compute, by zeroing real and imaginary parts, the zeros closer to $t=0$ of what here are called X1 and X3, the result is: $\pm 9.443$, $\pm 8.457$, $\pm 6.184,\ \ \pm 4.133\$. Very comparable with results in figure.
Figure 4. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=5$, with $p_{max}=(p^{*})^{3}=45.8\times 10^{6}$ ( here are used the first 2763823 primes in (3.16 ) ).The peaks are lower and fatter with respect to fig. 3, but, the inter-peaks behavior is much more close to $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$, or, better, to (4.4), that follows different behavior close to $t=0$ if $\alpha=0$ (even chrs) or $\alpha=1$ (odd chrs).
Figure 5. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=3$ and $q=4$, with $p_{max}=(p^{*})^{3}=158\times 10^{6}$ ( here are used the first 8868881 primes).For $q=3$, and, $\chi_{1}(n)=(0,1,-1)$ or $\angle[\chi_{1}(n)]=(undefined,0,\pi)$ to comply with format of tab. 1. For $q=4$, and $\chi_{1}(n)=(0,1,0,-1)$, or, $\angle[\chi_{1}(n)]=(undefined,0,undefined,\pi)$ to comply with format of tab. 1. Notice close to $t=0$ the characteristic behavior of odd primitive L-functions. Compare with fig. 2. Until now the only example of even primitive L-function is phase variation of X2Q5 in fig. 3, and fig. 4. In this case compare the inter-peaks level with fig. 15 and asymptotic formula (A.14 ).
@@
Table 4. Characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=9$, $w=e^{i\pi/3}$. The table is loaded from [12] .
Figure 6. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab. 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes ).
Figure 7. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab. 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes).
Figure 8. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) flipped vertically. So Lemma 2 is verified.
Figure 9. The expression (3.16 ) in character case $q=13$, and, for $\chi_{10}$ has been computed with $\epsilon=0$ and $\epsilon=0.1$. Notice that with greater $\epsilon$ the curve is close to $t$ axis as predicted in Lemma 6 . See (3.26 ). The choices are: $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).
Figure 10. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes). Notice that inter-peaks level is the same of fig, (15) flipped vertically. So Lemma 2 is verified.
Figure 11. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 12. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 13. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (15) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 14. The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (15) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.